OpenStudy (anonymous):

Determine which ordered pair is a solution of y = 8x A.(0,8) B.(-1,8) C.(1.5,10) D.(2,16)

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Nnesha

OpenStudy (texaschic101):

this is just a matter of subbing in your answer choices into the equation to see if it is true. I will do the first one for you... y = 8x using (0,8)...x = 0 and y = 8 now we sub 8 = 8(0) 8 = 0 (this is not a solution because it is false you wanna ry the next answer choice ?

OpenStudy (texaschic101):

* try

OpenStudy (anonymous):

Let me try the next one. Kind of see if I understand :)

OpenStudy (texaschic101):

take your time...I am not leaving

OpenStudy (anonymous):

8x, -1 = x and 8 = y 8 = 8(-1) 8 = -8, it isnt true. So B isnt an option

OpenStudy (anonymous):

Am I correct? :)

OpenStudy (texaschic101):

you definitely are correct...way to go. wanna try the next one ?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

x = 1.5, y = 10 8 = 1.5(10) 8 = 15 is not correct, That leaves D. But I want to make sure it is D.

OpenStudy (anonymous):

x = 2, y = 16 8 = 2(16) 2(16) = 32.........Not the answer....Now I am confused o.O

OpenStudy (texaschic101):

you got your numbers (x and y) mixed up y = 8x (1.5,10) 10 = 8(1.5) 10 = 12 (it is still false, but be careful of where you put your x and y numbers

OpenStudy (anonymous):

Oooooooh. Oops. Sorry! Lol

OpenStudy (texaschic101):

y = 8x.....(2,16)...x = 2 and y = 16 16 = 8(2) 16 = 16 (correct....so it is D)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Not really. Still kinda confused tho. I dunno why, because you stated how to do you clearly.

OpenStudy (texaschic101):

just remember, in all points, the x is the first number and the y is the second (x,y)

OpenStudy (anonymous):

Alright. Thank you! :)

OpenStudy (texaschic101):

anytime :)